OpenStudy (anonymous):

Divide: 4x-8/x^2-3x+2 divide 3x-6/x-1

OpenStudy (anonymous):

Factored out: 4(x+2)/(x+2)(x-1) divide 3(x+2)/x-1

OpenStudy (anonymous):

\[\frac{(4x-8)/(x{}^{\wedge}2-3x+2)}{(3x-6)/(x-1)}=\frac{(4 (-2+x))/((-2+x) (-1+x))}{(3 (-2+x))/(x-1)} \]Did you get \[\frac{4}{3 (x-2)} \]

OpenStudy (anonymous):

It's setup like this: \[\frac{ 4x-8 }{ x^{2}-3x+2 } \div \frac{ 3x-6 }{ x-1 }\]

OpenStudy (mertsj):

Remember that you have to multiply by the reciprocal of the division. So change the sign to multiplication and tip the second fraction upside down.

OpenStudy (mertsj):

Then factor everything that can be factored and cancel the common factors.

OpenStudy (anonymous):

Before I do that, was the factoring correct? (2nd post) @Mertsj

OpenStudy (anonymous):

 ?